OpenStudy (anonymous):

Can someone explain to me how to solve 4y=7x+7?

OpenStudy (anonymous):

as there are two unknowns so u need two equations to solve

OpenStudy (anonymous):

sorry, i forgot the other number is 112

OpenStudy (anonymous):

4y+112=180 y=68/4 y=17 7x+7= 112 x+1=112/7 x+1=16 x=15

OpenStudy (anonymous):

first supplementary angles then vertically opposite angle

OpenStudy (anonymous):

wow ok i see, i just misunderstood the problem completely. thank you